# FutureGen: LLM-RAG Approach to Generate the Future Work of Scientific Article

### v1

@@ 1 Introduction
Author-written future work sections have many problems. Firstly, they are often unspecific, ambiguous, not always easy to find, and speculative in nature Suray et al. (2024). Secondly, authors may struggle to articulate meaningful future directions, particularly when faced with space constraints or time limitations. Thirdly, they may hesitate to share research plans without a clear reward. This may partially explain why many future work proposals are overlooked post-publication Teufel (2017).
@@
  • Dataset. We create a dataset of future work sentences from nearly 8,000 papers collected from the ACL conference from 2012 to 2024. Papers often do not have exclusive future work sections (they are combined with general conclusions or limitations), so we use LLM to extract appropriate sentences. We then validated this extraction process with human annotators, ensuring that the LLM accurately identified future-work content.
@@
Figure 1: Overview of the future work Generation System using LLM and RAG.
@@ 3 Dataset Collection and Extraction
We extracted future work sentences from approximately $8,000$ papers published in the ACL conference from 2012 to 2024 using the Science Parse [^1] https://github.com/allenai/science-parse tool. The extraction process is as follows: (1) If a paper has a section explicitly labeled “Limitations and future work”, we extract the entire section. (2) If “future work” is not in the section title, we extract sentences with at least one of the phrases “future” or “future work”. Such regex-based string matching has a high recall, and we further filter these sentences using LLMs to improve precision (Details  5.2) (§4). This produces future work paragraphs from 6,227 papers, averaging five sentences per paper with an average word length of 65.
@@ 4 Methodology
Our methodology consists of two steps:
(1) Generating Topics and Titles of Future Work: We applied Topic Modeling with LLM to generate topics and titles for trend analysis of future work Al Azher et al. (2024).
(2) Generating future work: We used an LLM-based RAG approach to generate future work from the key sections of research papers with LLM-based evaluation and feedback.
@@ 4.1 Generating Topics and Titles of future work
Our dataset spans 13 years, with each year containing $n$ research papers. We extracted its future work section, denoted as $FW={fw_{1},fw_{2},....fw_{n}}$, where $fw_{i}$ represents the ‘future work’ paragraph containing a bunch of sentences from paper $p_{i}$. We then input these texts into BERTopic, which clusters them into $x$ topics (e.g., $t_{1},t_{2}...t_{x}$). Each topic $t_{i}$ comprises a set of topic words $tw_{1},tw_{2}...tw_{z}$ and a corresponding collection of relevant texts $R_{i}$. We utilized a topic modeling approach named BERTopic, utilized a transformer-based embedding from BERT convert to dense vector, and then created clusters to identify groups of similar documents. To enhance interpretability, we integrate an LLM) with BERTopic to generate a descriptive title for each topic (see Figure 3 in the Appendix). This entire process is repeated for every year in our dataset and collected topic title for each year (Table 9, 7, Appendix).
@@ 4.2 Generating Future Work
Figure 1 illustrates the workflow of our proposed model. We choose $X$ random papers from the dataset and store them into a vector database. Later, we select the top three sections most similar to "future work" using cosine similarity (Task 1) (step 1a) and send them to the Retriever, which gathers relevant information from a vector database (steps 2a-5). The augmented input is then processed by the LLM generator to generate future work (steps 6,7) (Task 3). Simultaneously, extracted future work from a paper is refined (re-extracted) by the LLM to remove sentences unrelated to future work (steps 1b,2b) and make ground truth (Task 2). Here, the ground truth of each paper contains a bunch of future work as a paragraph. Later, LLM-generated future work and ground truth are evaluated by the LLM (step 8, Figure 1), which assigns scores based on performance (step 9). If the generated output does not meet the threshold, the LLM re-generates it based on feedback (steps 10, 11) (Task 4).
@@
Task 1: Section Selection. After extracting the texts (Section 3), we used a cosine similarity-based approach to identify the most relevant sections for generating future work. Each paper consists of sections such as Abstract, Introduction, and Methodology. We calculated the cosine similarity between each section and the paper’s future work section (Table 8, Appendix) across all papers, and we selected the top three sections with the highest average similarity, which were Abstract, Introduction, and Conclusion. These sections formed the basis for generating future work content.
Task 2: Refining Ground Truth. After extracting the future work sections described in Section 3, we employed an LLM as an extractor role to isolate only future work sentences while removing irrelevant sentences, resulting in a Refined future work (Figure 1, step 1b). Extracting the future work is challenging since most papers do not have a dedicated section, and it is often scattered in any other section. Therefore, filtering out noisy sentences is crucial to ensure accurate extraction ( 5.2). Additionally, we conducted a human evaluation to assess whether LLM is good at extracting future work sentences or not. By this process, we made silver standard data (Details in  6.4).
@@
Task 3: Generating future work Using LLMs and RAG. Generating future work from a single paper may lead to narrow, redundant, or overly specific suggestions, missing broader research trends and interdisciplinary insights. Incorporating other papers’ content through RAG enhances contextual relevance, ensuring alignment with ongoing advancements and diverse, meaningful contributions reflective of real-world research priorities. To generate the future work of a paper, we utilized the top three sections (Abstract, Introduction, and Conclusion) selected in Task 1, integrating a RAG system, which consists of ‘X’ number of research papers (Details  5.3). The Retriever processes the input query, which includes the prompt and the top three sections of the research paper. It then retrieves additional relevant information from the vector database using cosine similarity and applies a ranking mechanism to prioritize the extracted text (Figure 1, steps 2a,3a,4,5). The final augmented input (a combination of the prompt, selected sections, and ranked retrieved text)is fed into the LLM Generator, producing future work content based on the provided context (Figure 1, step 6).
@@
Task 4: Evaluation and Iterative Refinement. After generating future work from Task 3, we evaluated its quality against the Refined future work (ground truth from Task 2). We employed several NLP-based quantitative metrics and used an LLM as a Judge to assess LLM-generated text against the Refined future work based on coherence, relevance, readability, grammar, and overall quality (Figure 1, step 8,9). Each evaluation was rated on a 1 (worst) to 5 (best) scale, with justifications provided (Figure 4, Appendix). We set a threshold of 3 as an acceptable midpoint. If an LLM-generated future work received a score of 3 or lower in any metric, the justification was incorporated into the prompt, and the future work was regenerated accordingly. This iterative refinement process was repeated up to two times to assess whether performance improved (Figure 1, step 9,10 11).
@@
ROUGE-1 24.33 25.74(
+1.41
ROUGE-2 5.27 7.85(
+1.05
ROUGE-L 17.24 20.25(
+2.58
BScore(f1) 87.23 87.50(
+0.27
Jaccard S 15.40 18.13(
+2.73
Cosine S 48.07 57.03(
+8.96
BLEU 2.38 5.74(
+3.36
Coherence 3.94 3.97(
+0.03
Relevance 4.07 4.62 (
+0.55
Readability 3.19 3.49 (
+0.30
Grammar 4.02 4.01 (
-0.01
Overall 3.85 3.95 (
+0.10
Table 1: Comparison of performance without and with LLM feedback using GPT-4o mini. Metrics include Jaccard Similarity (Jaccard S) and Cosine Similarity (Cosine S).
@@
Table 2: Performance of different models in generating future work, evaluated against paper ground truth extracted by LLM (GPT Refined Ground Truth). Note: GPT 4om refers to GPT-4o mini, ZS indicates Zero-Shot, FT represents Fine-Tuning, It. denotes Iteration, Coh. stands for Coherence and Logic, Rel. for Relevance and Accuracy, Read. for Readability and Style, and Gr. for Grammatical Correctness.
Model Sec. Iter. R-1 R-2 R-L BS-F1 JS CS BLEU
RAW Ground Truth
@@
GPT 4o mini 3 1 17.50 3.11 14.22 85.91 11.55 40.53 1.29
GPT 4o mini + RAG 3 1 21.20 5.09 16.41 86.40 13.27 40.36 2.84
GPT Refined Ground Truth
@@
GPT 4o mini 3 1 24.33 5.27 17.24 87.23 15.40 48.07 2.38
Llama 3 ZS 3 1 14.15 3.23 10.38 85.61 14.52 52.27 2.27
Llama 3 FT 3 1 26.75 5.61 - 85.54 12.28 45.37 1.14
GPT 4o mini 3 1 24.59 5.62 17.69 87.44 15.87 50.52 2.72
GPT 4o mini + RAG 3 1 21.13 5.52 14.19 87.67 16.95 49.32 3.42
GPT 4o mini + RAG 3 2 26.58 7.97 20.87 88.15 18.67 58.33 5.67
GPT 4o mini + RAG 3 3 26.96 7.53 20.18 87.94 17.92 56.34 4.68
@@
Table 3: Performance comparison of various models in generating future work, considering Extracted future work from papers using Python regex (RAW Ground Truth, RAW GT) and LLM-refined future work extracted from RAW GT (GPT Refined Ground Truth, GPT GT). Note: Iter. denotes Iteration, Sec. refers to Section, R-1, R-2, and R-L represent ROUGE-1, ROUGE-2, and ROUGE-L scores, respectively. BS-F1 indicates BERTScore (F1), while JS and CS stand for Jaccard Similarity and Cosine Similarity, respectively.
5 Experimental Setup
5.1 Topic Modeling
For trend analysis, we employed topic modeling enhanced with an LLM to generate both topics and corresponding titles. Specifically, we integrated GPT 4o mini in a zero-shot configuration with BERTopic (since few-shot approaches tend to overfit by mirroring prompt examples) and set a minimum similarity threshold of 0.75.
5.2 Extraction
We employed the Science Parse tool and python regex to extract each paper’s future work sections. After isolating these texts, we applied GPT 4o mini with a temperature of 0 in a zero-shot configuration. Our prompt, “Extract sentences related to future work, nothing else,” successfully guided the model to extract a coherent paragraph comprised only of future work-related sentences, effectively discarding unrelated content.
5.3 Generation
To generate future work texts, we fine-tuned BART and T5 models using 70% training data and 30% testing data. BART has a maximum token limit of 1024, while T5 is limited to 512 tokens, and any input exceeding the respective token limits is disregarded. For both models, the generated text is capped at 100 tokens. Additionally, we applied GPT-3.5 and GPT-4o using a zero-shot approach with a set temperature of 0.2 for the same tasks. The prompts mentioned that the generated text shouldn’t be more than 100 words and leveraged their 16,000 and 128,000 token context windows, respectively.
Generating Cosine Similarity: We used Sentence Transformers (’all-MiniLM-L6-v2’) for embedding generation and scikit-learn’s  [^2] https://scikit-learn.org/stable/ cosine similarity function for similarity computation.
Fine-Tuning LLaMA: For LLaMA 3.1 fine-tuning, it was trained using formatted data, where ‘Abstract,’ ‘Introduction,’ and ‘Conclusion’ served as input, and extracted ‘future work’ was used as output. LLaMA already contains extensive general knowledge, so they require few domain-specific examples to adapt effectively. So, we split the dataset into 30% training and 70% testing. We applied LoRA Hu et al. (2022) with 4-bit quantization (QLoRA) Dettmers et al. (2023) for memory optimization, in LoRA, we set alpha=16 without dropout as default parameter. FlashAttention Dao et al. (2022) was incorporated to improve processing speed. We set a maximum input sequence length of 2048 tokens and trained the model for 60 steps with a default learning rate of 2e-4. During testing, we provided a prompt with detailed instructions for generating future work, with a maximum output length of 128 tokens and a temperature setting 1.
LLM with RAG Integration We integrated RAG with GPT-4o mini, leveraging OpenAI’s ‘text-embedding-3-small’ model. Our vector database comprises 100 randomly selected papers from the dataset, and these papers were removed from the dataset. For semantic search using vector embeddings, we employed LlamaIndex, utilizing its in-memory vector store rather than an external database. We segmented the data into chunks of up to 512 tokens to accommodate smaller context windows. The overall context window was set to 3,900 tokens—the maximum number of tokens the LLM can process at a time. Additionally, we set $K=3$, meaning that the top three most relevant chunks are retrieved from the vector store to provide contextual support during the generation process.
For LLM as a judge and LLM as an extractor, we used a one-shot approach with GPT 4o mini and a zero-shot approach with GPT 4o mini, respectively. We experimented with GPT 4o mini vs Llama 3-80b as a Judge and found Llama 3-80b provides a higher score than GPT 4o mini (11, Appendix).
6 Experiments and Results
6.1 Trend Analysis of future work through Topic Modeling
For trend analysis of future work, we chose BERTopic, and after testing various sentence transformer models, we selected ‘all-MiniLM-L6-v2’ (Table 16, Appendix). Additionally, we tested KMeans clustering with 15-30, but in all cases, silhouette and coherence scores dropped. We also experimented with UMAP, varying neighbors, and components (5–20), but its integration deteriorated performance (Figure 3, Appendix).
6.2 Evaluation of Extraction
Evaluation-raw ground truth vs. GPT-refined ground truth. We experimented with two ground truth approaches: (1) the raw ground truth (RAW GT) and (2) the GPT-refined ground truth (GPT GT) with LLM-generated text. RAW GT refers to the future work texts extracted from research papers using ScienceParse tool and regular expressions in Python. GPT GT takes RAW GT and utilizes the LLM to extract only sentences related to future work. After generating future work from key sections using our proposed approach, we evaluated ‘generated future work’ with RAW GT and GPT GT. As shown in (Table 6, Appendix), GPT GT outperformed RAW GT across all metrics, demonstrating improved text quality and relevance when using GPT 4o mini. Additionally, Table 3 confirms that GPT GT yields better performance than RAW GT across different models when comparing the same model’s RAW GT with GPT GT.
6.3 Evaluation of Generation
We experimented with various LLMs (Table 3) to generate future work sections. Traditional NLP-based metrics primarily focused on lexical overlap rather than semantic quality, while LLM-based evaluation provides a more contextual assessment than token-level comparison. Alongside the NLP-based evaluation, it focused on coherence and logic, relevance and accuracy, readability and style, grammatical correctness, and overall impression.
Evaluation-considering all sections vs top-3 sections. Instead of processing the entire paper, we used cosine similarity to select the three most relevant sections as input. As shown in (Table 4, Table 3), using only the top three sections resulted in a slight performance drop (e.g., ROUGE-L: -0.38, BERTScore: -0.12).
Metrics All sec 3 sec.
@@ 6.3 Evaluation of Generation
Table 4: Performance comparison between using three selected sections versus all sections as input for future work generation with GPT-3.5

### v2

@@ 1 Introduction
Author-written future work sections have many problems. Firstly, they are often unspecific, ambiguous, not always easy to find, and speculative in nature Suray et al. (2024). Secondly, authors may struggle to articulate meaningful future directions, particularly when faced with space constraints or time limitations. Thirdly, they may hesitate to share research plans without a clear reward. This may partially explain why many future work proposals are overlooked post-publication Teufel (2017). To alleviate this problem, we generate future work from each paper and evaluate it against both the author’s mentioned future work and long-term goals extracted from OpenReview peer reviews in our approach. By merging these peer-reviewed objectives with the author-mentioned future work, we build a more comprehensive and robust ground truth.
@@
  • Dataset. We created a dataset of future work sentences from nearly 8,000 papers collected from the ACL conference from 2012 to 2024 and 1000 papers from NeurIPS. Papers often do not have exclusive Future Work sections (they are combined with general conclusions or limitations), therefore, we use LLMs to extract relevant sentences. We also validated this extraction process on a random sample of the dataset with human annotators, ensuring that the LLM accurately identifies future-work content.
@@
Table 1: Future-Work Suggestions from Different Sources
Source Future-Work
Ground truth We will explore the application of SWAD to other robustness problems, including ImageNet classification and its robustness benchmarks.
LLM Future research could explore the applicability of flatness-aware methods in other machine learning contexts, such as reinforcement learning or unsupervised learning.
RAG Investigate the effectiveness of the SWAD method across a wider variety of datasets and domains, particularly those with significant distribution shifts. This would validate the generalization capabilities of SWAD beyond the current benchmarks.
@@ 3 Dataset Collection and Extraction
We created our dataset by extracting Future Work sentences from approximately $8,000$ papers published in the ACL conference from 2012 to 2024. We also collected $1,000$ papers from NeurIPS, along with their open-access peer reviews from OpenReview  [^1] https://openreview.net from 2021-22. The extraction process from the paper is as follows: (1) If a paper has a section explicitly labeled “Limitations and Future Work”, we extract the entire section using the Science Parse [^2] https://github.com/allenai/science-parse tool (2) If “Future Work” is not in the section title, we extracted sentences with at least one of the phrases “future” or “Future Work" and collect all of the sentences up to the next section. We employed python regex for this purpose. Such regex-based string matching has a high recall, so we further filtered these sentences using LLMs to improve precision. Since most papers do not have a dedicated section, it is often scattered in any other section. Therefore, filtering out noisy sentences is crucial to ensure accurate extraction. Here we employed an LLM as an extractor role to isolate only Future Work sentences while removing irrelevant sentences. This produces Future Work paragraphs from 6,227 papers from ACL and 1000 papers fron NeurIPS, averaging five sentences per paper with an average word length of 65.
OpenReview: Since no ACL OpenReview were available during our data collection process, we collected OpenReview from the NeurIPS papers only. After parsing text from OpenReview, we gathered all peer feedback and used an LLM to extract potential future-work suggestions for the authors. We then applied a second LLM to validate each extracted sentence, discarding any that did not represent true long-term research goals.
3.1 Human Evaluation
Additionally, we conducted a human evaluation to assess whether LLM is good at extracting Future Work sentences or not. We conducted a user study with 200 randomly selected samples involving three annotators who rated below two questions from 1 to 3 (1 = worst, 3 = best) with the question “How well does LLM extract the Future Work (ground truth) without generating/hallucinating? (Q1)”. The annotators are graduate students, knowledgeable about machine learning, and not related with this paper. We observed a mean user rating of 2.54 out of 3 (Table 11), indicating strong agreement; this procedure thus produced our silver-standard dataset.
@@ 4 Methodology
Our workflow consists of two main stages: 1. Generating Topics and Titles of Future Work: We applied Topic Modeling with LLM to generate topics and titles for trend analysis of future work Al Azher et al. (2024).
2. Generating Future Work Text: We used an LLM-based RAG approach to generate future work from the key sections of research papers with LLM-based evaluation and feedback.
@@ 4.1 Generating Topics and Titles of future work
Our first work focuses on generating topics and titles from author-stated future work for trend prediction from ACL data. We applied Topic Modeling with LLM to generate topics and titles for trend analysis of Future Work (Figure 1). Here, the input consists of author-stated future work from various papers published in a single year, and the output is a set of generated titles.
Figure 1: BERTopic + LLM pipeline for generating titles from Future Work.
Our dataset spans 13 years, containing approximately $n$ research papers in each year. We extracted the author’s mentioned Future Work from $n$ research papers, denoted as $FW={fw_{1},fw_{2},...,fw_{n}}$, where $fw_{i}$ represents the ‘Future Work’ paragraph containing sentences from paper $p_{i}$. These input texts are sent into BERTopic, a topic modeling tool that utilizes a transformer-based embedding from BERT. BERTopic clustered them into $x$ topics (e.g., $t_{1},t_{2}...t_{x}$). Each topic $t_{i}$ comprises a set of topic words $tw_{1},tw_{2}...tw_{z}$ and a corresponding collection of relevant texts $R_{i}$. To enhance interpretability, we integrate an LLM with BERTopic to generate a descriptive title for each topic (Figure 1). We tuned the components of BERTopic with UMAP, Count Vectorizer, Sentence Transformer, HDBSCAN, and Zero-shot min similarity for guided topic modeling. This entire process is repeated for all articles in each year in our dataset and the number of topic titles collected $x$ for each year is presented in (Table 15).
@@ 4.2 Generating Future Work
The second work focuses on an LLM-based RAG pipeline that generates future-work sections directly from ACL and NeurIPS papers and refines them through LLM-driven evaluation and feedback. (Figure 2). In the ACL dataset, the ground truth consists of author-stated future work statements extracted using an LLM. For the NeurIPS dataset, the ground truth includes both the authors’ future work statements and long-term suggestions from OpenReview peer reviewers. In both cases, we constructed input data by collecting all texts from the full paper after removing ‘author-mentioned future work’. The output is the future work text generated by the model.
Figure 2: Overview of our LLM + RAG future-work pipeline: we extract author-written future work from the paper, extract long-term goals from peer reviews via a tool + LLM, generate new suggestions with a RAG-augmented LLM from the paper, evaluate their quality, and apply self-feedback for refinement.
Figure 2 depicts our end-to-end pipeline: starting from a sample paper, we first extract candidate “future work” sentences via regex and refine them with an LLM, then pull peer-review comments from OpenReview and use the same LLM to isolate long-term goals—merging both into a robust ground truth. We remove the author’s original future work, feed the paper to an LLM augmented by a vector-store retriever that supplies related documents, and generate new future work suggestions. Each paper is automatically scored on NLP metrics and LLM-based criteria (coherence, relevance, novelty, grammar, overall impression), and any suggestion scoring below a threshold of 3 is critiqued and re-fed into the generator for a polished second iteration.
A more detailed overview of the tasks in this stage is given below.
@@
Task 1: Section Selection. After extracting the texts (Section 3), we used a cosine similarity-based approach to identify the most relevant sections for generating Future Work. We calculated the cosine similarity between each section and the paper’s Future Work text (Table 14, Appendix) across all papers, and we selected the top three sections with the highest average similarity. These sections formed the basis for generating Future Work content. For all experiments, we removed the author’s mentioned future work from the paper and made the input text.
@@
Task 2: Generating Future Work Using LLMs and RAG. Generating Future Work from a single paper may lead to narrow, redundant, or overly specific suggestions, missing broader research trends and interdisciplinary insights. For example, a zero-shot LLM often defaults to boilerplate suggestions like “explore more data” or “apply to new domains.” RAG combats this by feeding the model real, paper-specific passages at generation time, anchoring its outputs in concrete evidence and substantially reducing hallucinations. Moreover, by incorporating content from related papers, RAG ensures that future-work sections remain contextually relevant, tailored to genuine research priorities. To generate the Future Work of a paper, we experimented with the top three sections selected in Task 1, and all sections (excluding future work) integrating a RAG system, which consists of 100 research papers (Details  5.1). The Retriever processes the input query, which includes the prompt and the content of the research paper. It then retrieves additional relevant information from the vector database using cosine similarity and applies a ranking mechanism to prioritize the extracted text (Figure 2). The final augmented input is fed into the LLM Generator, producing Future Work content based on the provided context (Figure 2, step 6).
@@
Task 3: Evaluation and Iterative Refinement. We employed several NLP-based quantitative metrics and used an LLM as a Judge to assess LLM-generated text against the ground truth Future Work based on coherence, relevance, readability, grammar, overall quality, and novelty (Figure 2, steps 8 and 9). Each evaluation was rated on a discrete scale of 1 (worst) to 5 (best), with justifications provided except novelty which is 1-10. We set a threshold of 3 as an acceptable midpoint, 7 for novelty. If an LLM-generated Future Work received a score less than or equal to the midpoint in any metric, the justification was incorporated into the prompt, and the Future Work was regenerated accordingly. This iterative refinement process was repeated up to two times to assess whether performance improved (Figure 2, step 9, 10, and 11).
Meth. Abs. Intro. RW Data Meth. Exp. Con. Lim.
CS 25.08 24.52 21.09 8.96 14.82 21.79 22.0 21.38
Table 2: Average Cosine Similarity of each section with Future Work in ACL data. N.B: CS, Abs, Intro, RW, Data, Meth, Exp, Con, Lim means Cosine Similarity, Abstract, Introduction, Releated Work, Data, Methodology, Experiment, Conclusion, and Limitations, respectively.
@@
ROUGE-1 24.33 25.74( +1.41 )
ROUGE-2 5.27 7.85( +1.05 )
ROUGE-L 17.24 20.25( +2.58 )
BScore(f1) 87.23 87.50( +0.27 )
Jaccard S 15.40 18.13( +2.73 )
Cosine S 48.07 57.03( +8.96 )
BLEU 2.38 5.74( +3.36 )
Coherence 3.94 3.97( +0.03 )
Relevance 4.07 4.62 ( +0.55 )
Readability 3.19 3.49 ( +0.30 )
Grammar 4.02 4.01 ( -0.01 )
Overall 3.85 3.95 ( +0.10 )
Table 3: Comparison of performance without and with LLM feedback using GPT-4o mini in ACL data. Metrics include Jaccard Similarity (Jaccard S) and Cosine Similarity (Cosine S).
GT Iter. R-1 R-2 R-L BS CS JS Bl CST JST Coh Rel Read Gram Nov Ov
GPT 4o mini
FW 1 26.03 10.28 15.54 85.02 68.23 11.21 2.29 33.02 74.11 4.31 4.63 3.82 4.81 7.28 4.29
OR 1 26.03 6.97 12.34 82.94 62.56 10.43 1.26 18.75 62.07 3.67 4.37 3.19 4.18 7.84 3.65
FW+OR 1 37.98 12.94 18.32 85.20 74.33 14.46 3.55 33.67 76.78 4.41 4.87 3.85 4.67 7.28 4.40
FW+OR 2 37.35 12.14 17.43 84.63 73.50 14.29 3.42 31.58 68.42 4.48 4.79 4.15 4.89 8.00 4.44
GPT 4o mini + RAG
FW 1 21.79 6.14 11.89 83.00 63.27 8.72 1.08 21.01 62.79 4.26 4.04 4.87 5.00 7.70 4.15
OR 1 26.12 6.72 12.43 82.81 62.91 10.15 1.01 18.54 51.67 4.56 4.56 4.89 5.00 7.80 4.52
FW+OR 1 34.28 9.57 15.90 83.47 72.57 12.29 2.02 24.16 62.97 4.17 4.36 4.89 5.00 7.11 4.38
Table 4: Performance comparison between different methods in NeurIPS data. Here, GT, FW, OR, Iter means Ground Truth, Author Mentioned Future Work, OpenReview suggestion, and Iterations, respectively. Metrics include ROUGE (R-1, R-2, R-L), BERTScore (BS), Cosine Similarity (CS), Jaccard Similarity (JS), BLEU (Bl), Coherence (Coh), Relevance (Rel), Readability (Read), Grammar (Gram), Novelty (Nov), Overall (Ov), and Similarity (Sim).
@@
Table 5: Performance of different models in generating Future Work, evaluated against paper ground truth extracted by LLM (GPT extracted Ground Truth) in ACL data. Note: GPT 4om refers to GPT-4o mini, ZS indicates Zero-Shot, FT represents Fine-Tuning, It. denotes Iteration, Coh. stands for Coherence and Logic, Rel. for Relevance and Accuracy, Read. for Readability and Style, and Gr. for Grammatical Correctness.
5 Experimental Setup
5.1 Generating Topics and Titles of Future Work
Our first work, trend analysis, we leverage GPT-4o-mini’s deep semantic understanding to produce concise, human-readable topic titles on top of BERTopic’s clustering. We set the minimum similarity threshold to 0.75 to ensure that each topic cluster remains semantically tight (boosting coherence) without collapsing into overly narrow groups. After benchmarking several sentence-transformer embeddings, we selected all-MiniLM-L6-v2 for its superior coherence. Alternative setups—including KMeans clustering (15–30 clusters) and UMAP dimensionality reduction (5–20 neighbors/components)—consistently produced lower silhouette and topic coherence scores (Figure 1). This LLM-augmented, prompt-light setup gives us automatically generated, high-quality topic labels at scale, with no manual annotation or prompt-tuning overhead.
Our second work is future work generation, where we benchmarked a diverse mix of generative and retrieval-augmented methods to understand their relative strengths under realistic constraints. BART and T5 serve as strong, well-studied seq2seq baselines with fixed token-limit trade-offs, while GPT-3.5 and GPT-4o illustrate how off-the-shelf LLMs perform zero-shot under a controlled prompt budget. Fine-tuning LLaMA-3.1 with LoRA/QLoRA and FlashAttention demonstrates that even large open-source models can be adapted efficiently on modest hardware. Finally, integrating RAG grounds generation in concrete evidence, and comparing one-shot versus zero-shot LLM evaluators (GPT vs. Llama) lets us quantify both generation quality and evaluator bias. This multi-axis evaluation ensures our conclusions generalize beyond any single model or configuration.
Generating Similarity between Other Sections and Future Works: We used Sentence Transformers (’all-MiniLM-L6-v2’) for embedding generation and scikit-learn’s  [^3] https://scikit-learn.org/stable/ cosine similarity function for similarity computation.
Fine-Tuning Models: First, we fine-tuned BART (1,024-token limit) and T5 (512-token limit) on a 70/30 train/test split, discarding any over-length inputs. For LLaMA 3.1 7B fine-tuning, it was trained using ‘Abstract,’ ‘Introduction,’ and ‘Conclusion’ as input, and the extracted ‘ Future Work’ was used as output. Leveraging LLaMA’s extensive pretraining, we fine-tuned it with QLoRA (4-bit, alpha=16) and FlashAttention on a 30/70 train/test split for 60 steps (learning rate 2e-4, 2048-token context). During testing, we provided a prompt with detailed instructions for generating Future Work, with a maximum output length of 128 tokens and a temperature setting of 1.
Zero/Few Shot(s) LLM: We ran GPT-3.5 and GPT-4o in zero-shot mode, leveraging their 16 K and 128 K context windows, respectively. For LLM as a judge and LLM as an extractor, we used a one-shot approach with GPT-4o mini and a zero-shot approach with GPT-4o mini, respectively.
LLM with RAG Integration: We integrated RAG with GPT-4o mini for Future work generation, leveraging OpenAI’s ‘text-embedding-3-small’ model for relevant document retrieval. Our vector database comprises 100 randomly selected papers from the dataset, and these papers were removed from the dataset. For semantic search using vector embeddings, we employed LlamaIndex, utilizing its in-memory vector store rather than an external database. We used a hybrid retriever system to fetch the data from the vector database consisting of BME and FAISS with an equal weight of 50%. We segmented the data into chunks of up to 512 tokens to accommodate smaller context windows. The overall context window was set to 3,900 tokens—the maximum number of tokens the LLM can process at a time. Additionally, we set $K=3$, meaning that the top three most relevant chunks are retrieved from the vector store to provide contextual support during the generation process.
6 Experiments and Results
Model Sec. Iter. R-1 R-2 R-L BS JS CS Bl
Tool extracted Ground Truth
@@ 6 Experiments and Results
GPT4om 3 1 17.50 3.11 14.22 85.91 11.55 40.53 1.29
GPT4om+R 3 1 21.20 5.09 16.41 86.40 13.27 40.36 2.84
GPT Extracted Ground Truth
@@
Llama 3 Z 3 1 14.15 3.23 10.38 85.61 14.52 52.27 2.27
Llama 3 F 3 1 26.75 5.61 - 85.54 12.28 45.37 1.14
GPT 4om 3 1 24.59 5.62 17.69 87.44 15.87 50.52 2.72
GPT4om+R 3 1 21.13 5.52 14.19 87.67 16.95 49.32 3.42
GPT4om+R 3 2 26.58 7.97 20.87 88.15 18.67 58.33 5.67
GPT4om+R 3 3 26.96 7.53 20.18 87.94 17.92 56.34 4.68
@@
Table 6: Performance comparison of various models in generating Future Work, considering Extracted Future Work from papers using Tool extracted Future Work and GPT-extracted Future Work in ACL data. Note: GPT 4om, R, Llama 3 Z, F, Iter. denotes GPT 4o mini, RAG, Llama 3 Zero shot, Fine Tuning, Iteration, Sec. refers to Section, R-1, R-2, R-L, BS, JS, CS, Bl represent ROUGE-1, ROUGE-2, ROUGE-L, BERTScore, Jaccard Similarity, Cosine Similarity, Bleu respectively.
Metrics All sec 3 sec. Tool extracted GT GPT extracted GT
@@
Table 7: Performance comparison of GPT-3.5 in generating Future Work using three selected sections versus full-text input, evaluated against both Tool extracted Ground Truth (GT) and GPT-extracted ground truth on ACL data.
6.1 Evaluation of Future Work Extraction
After extracting the author mentioned future work from a paper using a tool (Tool extracted), we send it to the LLM to re-extract the future work by removing noisy sentences (GPT Extracted). We measured NLP-based evaluation (Rouge, BERTScore, BLEU) and LLM-based evaluation (Coherence, Relevance, etc.) between LLM-generated text and Tool extracted Ground Truth (GT). We also evaluated LLM-generated text with GPT extracted GT. As shown in (Table 7), GPT extracted GT outperformed across all metrics, demonstrating improved text quality and relevance when using GPT 4o mini. Additionally, Table 6 confirms that GPT extracted GT yields better performance than Tool extracted GT across different models when comparing the same model’s Tool extracted GT. The reason is that Tool extracted GT often contains noisy sentences, and not all of the sentences are related to Future Work. Where GPT extracted GT conatins only sentences related to Future Work, achieves a high alignment with LLM generated text. We further conducted a human evaluation to evaluate the text extraction process by GPT in section 3.1.
6.2 Evaluation of Future Work Generation
We experimented with various LLMs (Table 6) to generate Future Work sections. Traditional NLP-based metrics primarily focused on lexical overlap and semantic quality, and LLM-based evaluation provides more contextual assessment and novelty. NLP-based metrics couldn’t measure novelty and relied more on ground truth, missing the depth evaluation. Solely relying on an LLM-based evaluation system raises potential bias issues. To alleviate these problems, we incorporated both LLM-based and NLP-based evaluation systems to make a robust evaluation system for evaluation by focusing on coherence and logic, relevance and accuracy, readability and style, grammatical correctness, overall impression, and novelty. Here, novelty measures how novel the generated future work is in terms of the ground truth (higher values indicate more original ideas).
6.3 Future Work Generation
Evaluation-considering all sections vs top-3 sections. We used cosine similarity to select the three most relevant sections as input in ACL data. We compared the performance considering 3 sections and all sections (excluding the ‘author’s mentioned future work’). As shown in Table 7 and Table 6, using the top three sections led to a slight performance drop in every metric, showing that the top 3 sections contained sufficient information to generate meaningful future work.